If the average of six numbers is 18 and the average of three of those numbers is 15, then what is the sum of the remaining three numbers? Must show all work.
Answer by ankor@dixie-net.com(22740)

If the average of six numbers is 18 and the average of three of those numbers is 15, then what is the sum of the remaining three numbers? Must show all work.
:
Find the sum of all 6 numbers with an average of 18
6 * 18 = 108
Find the sum of the 3 numbers with an average of 15
3 * 15 = 45
Find the sum of the remaining numbers
108 - 45 = 63